OpenStudy (diamondboy):

Find the limit as x approaches negative infinity of (6/x)-(x/4)

OpenStudy (diamondboy):

@freckles

OpenStudy (freckles):

hint: \[\frac{6}{x}-\frac{x}{4}=\frac{24-x^2}{4x} =\frac{\frac{24}{x}-\frac{x^2}{x}}{\frac{4x}{x}}=\frac{\frac{24}{x}-x}{4}\]

OpenStudy (diamondboy):

DNE?

OpenStudy (diamondboy):

or +ve infinity

OpenStudy (freckles):

While I do think you can say the limit dne, I think you could actually be more descriptive and say yes + inf since we would have \[\frac{0-(-\inf)}{4}=\frac{0+\inf}{4}=\frac{\inf}{4}=\inf\] And I hope you know what I just wrote is not formal at all like you can't actually plug in infinity as if it were a number :p

OpenStudy (diamondboy):

yep

OpenStudy (diamondboy):

I understand :)

OpenStudy (freckles):

cool stuff :)

OpenStudy (diamondboy):

how good are u with physics?

OpenStudy (diamondboy):

@freckles

OpenStudy (freckles):

not very good i actually had no physics class experience not even one class :p

OpenStudy (diamondboy):

k

OpenStudy (diamondboy):

I have one more question

OpenStudy (diamondboy):

Find the limit as x approaches - infinity of x/(SQRT(X^2-x))

OpenStudy (diamondboy):



OpenStudy (anonymous):

i think it will give infinity

OpenStudy (freckles):

you could divide both top and bottom by sqrt(x^2)

OpenStudy (freckles):

but you want to use that sqrt(x^2)=-x since x->-inf

OpenStudy (diamondboy):

dats -1?

OpenStudy (freckles):

\[\frac{\frac{x}{\sqrt{x^2}}}{\sqrt{1-\frac{1}{x}}}\]

OpenStudy (freckles):

actually yes you would have -1/sqrt(1-0) very nice

OpenStudy (diamondboy):

but can I do it like dis?